OpenStudy (anonymous):

HELP PLEASE!!! Solve 2x2 + 12x - 14 = 0 by completing the square. (2 points) 2(x + 6)2 = 26; x = -6 + , x = -6 - 2(x + 6)2 = 20; x = -6 + , x = -6 - 2(x + 3)2 = 32; x = 1, x = -7 2(x + 3)2 = 32; x = 13, x = 19

OpenStudy (anonymous):

do you have an idea of what the answer is?

OpenStudy (campbell_st):

wow... that's an interesting way to solve the equation... but rewrite the equation as \[2x^2 + 12x = 14\] factor out 2 on the left side \[2(x^2 + 6x) = 14\] now complete the square on left side....

OpenStudy (campbell_st):

I would have divided every term by 2 and you get \[x^2 + 6x - 7 = 0\] this equation can be fatored... find the factors of -7 that add to 6 the larger factor us positive (x + factor 1)(x - factor 2)